Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-18T12:00:47-0400

Lemma: Arbitrary union of open set is open set.

Proof:

Let fix "I=\{1,2,3...\}" and denote an open set "A_i" for each "i \in I" . Now we will show that

"U=\cup_{i=1}^{\infty} A_i"

is open. Consider any arbitrary "p \in U \implies \exist" at least one "A_j" such that "p \in A_j" which implies that there exist an open ball of radius "r_p" such that "B(p,r_p) \subset A_j" (Since, by our hypothesis every "A_i" is open "\forall i\in I" ,hence, "A_j" is open for some "i=j \in I" .)which implies that

"B(p,r_p) \subset \cup_{i=1}^{\infty}A_i \\
\implies B(p,r_p) \subset U"

Thus, "p" is interior point of "U" and as "p" is arbitrary, therefore all points of "U" are interior points .Hence "U" is open.


Let us denote "A=(a,b) \: \& S=[a,b]" open and closed intervals respectively.

Claim: "A" is open set in "\mathbb{R}"

Proof:

Let, "\forall q \in A" , "B(q,r_q)" is a open ball centered at "q" of radius "r_q >0" , Thus, clearly


"A= \cup_{q \in A} B(q,r_q)"

Therefore, from the above Lemma, implies "A" is open.


Now, consider "B'=\mathbb{R} \backslash B" is the complement of "B" ,thus if we prove that "B'" is open then "B" is a closed set in "\mathbb{R}" .

Clearly, "B'=(- \infty,a) \cup (b,\infty)" ,therefore again from the above Lemma "B'" is open (since, "(-\infty,a) \: \& (b,\infty)" are both open intervals).Hence, we are done.